Rasta jump from a building 34 m above the ground. What is the velocity of Rasta at the end of 2.5 seconds? What time ut takes to hit the ground? (Please show the solution)

1
Expert's answer
2021-11-16T13:41:12-0500

"v=gt=9.8\cdot2.5=24.5\ (s)"


"h=gt^2/2\to t=\sqrt{2h/g}=\sqrt{2\cdot 34/9.8}=2.63\ (s)"